# Huppert’s conjecture for finite simple exceptional groups of Lie type

### v1

@@ 6. Verifying Step 3: $G^{\prime}$ is isomorphic to $H$
Finally assume that $\sigma=\gamma\tau,$ where $\gamma$ is either trivial or a diagonal automorphism and $\tau$ is a graph automorphism. Let $\ell=\ell_{9f}(p)$ be a primitive prime divisor of $p^{9f}-1$ and let $s$ be a semisimple element of order $\ell$ in $(H^{*})^{\prime}.$ Let $T$ be a maximal torus of $H^{*}$ containing $s.$ It is well known that $T$ is uniquely determined up to conjugation in $H^{*}.$ We have that $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{9}$ and $\mathcal{G}^{F}$ possesses an irreducible character $\chi_{s}$ with degree $|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=|{\rm E}_{6}(q)_{ad}|_{p^{\prime}}/\Phi_{9},$ which is trivial on the center as $s\in(H^{*})^{\prime}$ so $\chi_{s}\in{\mathrm{Irr}}(H).$ As $\chi_{s}$ is invariant under the diagonal automorphisms, it suffices to show that $\chi_{s}$ is not $\tau$-invariant so it is not $\sigma$-invariant and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ which is impossible. Now we know that ${\mathbf{C}}_{H^{*}}(\tau)\cong\textrm{F}_{4}(q)$ and thus $\ell\nmid|{\mathbf{C}}_{H^{*}}(\tau)|.$ Consider the $H^{*}$-conjugacy class $(s).$ Since $|s|=\ell,$ we deduce that the $H^{*}$-conjugacy class $(s)$ is not $\tau$-invariant. By [22, Corollary 2.5], $\chi_{s}$ is not $\tau$-invariant. Thus $G=G^{\prime}C=G^{\prime}\times C,$ since $C\cap G^{\prime}=1,$ and hence $C\cong G/G^{\prime}$ is abelian.
The arguments for the remaining cases are similar. The proof of the main theorem is now complete.

### v3

@@ 7. Verifying Step 4: $G=G^{\prime}\times C_{G}(G^{\prime})$
Finally assume that $\sigma=\gamma\tau,$ where $\gamma$ is either trivial or a diagonal automorphism and $\tau$ is a graph automorphism. Let $\ell=\ell_{9f}(p)$ be a primitive prime divisor of $p^{9f}-1$ and let $s$ be a semisimple element of order $\ell$ in $(H^{*})^{\prime}.$ Let $T$ be a maximal torus of $H^{*}$ containing $s.$ It is well known that $T$ is uniquely determined up to conjugation in $H^{*}.$ Now $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{9}$ and $\mathcal{G}^{F}$ possesses an irreducible character $\chi_{s}$ with degree
$|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=|{\rm E}_{6}(q)_{ad}|_{p^{\prime}}/\Phi_{9},$
which is trivial on the center as $s\in(H^{*})^{\prime}$ so $\chi_{s}\in{\mathrm{Irr}}(H).$ As $\chi_{s}$ is invariant under the diagonal automorphisms, it suffices to show that $\chi_{s}$ is not $\tau$-invariant: if so, then it is not $\sigma$-invariant and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ which is impossible. We know that ${\mathbf{C}}_{H^{*}}(\tau)\cong\textrm{F}_{4}(q)$ and thus $\ell\nmid|{\mathbf{C}}_{H^{*}}(\tau)|.$ Consider the $H^{*}$-conjugacy class $(s).$ Since $|s|=\ell,$ $H^{*}$-conjugacy class $(s)$ is not $\tau$-invariant. By [26, Corollary 2.5], $\chi_{s}$ is not $\tau$-invariant.
The arguments for the remaining cases are similar. ∎
We now complete the proof of the main theorem.
Proof of Theorem 1.1.
Let $G$ be a finite group and let $H$ be a finite simple exceptional group of Lie type. In view of [12, 29, 30, 32, 33, 34], we assume that $H$ is one of the simple groups listed in (1). By Propositions 6.1 and 7.1, $G^{\prime}\cong H$ and $G\cong G^{\prime}\times C_{G}(G^{\prime})$. It follows that $C_{G}(G^{\prime})\cong G/G^{\prime}$ is abelian. Hence $G\cong G^{\prime}\times A$, where $A=C_{G}(G^{\prime})$ is an abelian group. This completes the proof of the theorem. ∎
Acknowledgment
The author is grateful to the reviewer and the editor for numerous suggestions and corrections that improve the exposition of the paper. The author also thanks Frank Lübeck for his help during the preparation of this paper.